# RetrySQL: text-to-SQL training with retry data for self-correcting query generation

### v1

@@ A.5. Training data example
Figure S2. Example of a training data sample used in our experiments. It consists of the following elements: DDL statements for schema representation, external knowledge and question extracted from the BIRD metadata, reasoning steps generated as described in Section 3.2 and the ground truth SQL query. These components are separated by special tokens to guide the model in the learning process.
A.6. Prompts
You are a SQL expert. When I provide you with a SQL query, your task is to describe step-by-step how a person would create such a query. Follow the standard SQL execution order. Write the steps from the perspective of someone constructing the query. If the query includes subqueries, describe them step-by-step in the same detailed manner as the main query before referencing them. Each step should represent a distinct operation. Make the operations as granular as possible.
Here is the standard SQL execution order you should follow for your explanation. If a given clause is not present in the query, skip it without mentioning its absence:
1. FROM clause (including JOINs).
2. WHERE clause.
3. GROUP BY clause.
4. HAVING clause.
5. SELECT clause.
6. ORDER BY clause.
7. LIMIT clause.
For each query I provide:
1. Explain the query step by step in plain language.
2. Ensure that each step corresponds to one small, logical operation.
3. Use clear and concise language for each operation.
4. Each step should be provided in a single line (use single newline character between steps).
Here is an example query for your reference:
SELECT T1.name, T1.email, SUM(T3.amount) AS total_sales FROM Customers AS T1 INNER JOIN Orders AS T2 ON T1.customer_id = T2.customer_id LEFT JOIN OrderDetails AS T3 ON T2.order_id = T3.order_id WHERE T2.order_date >= ’2023-01-01’ AND T2.order_date <= (SELECT order_date FROM Orders WHERE order_id = ’1’ ORDER BY order_date DESC LIMIT 1) GROUP BY T1.name, T1.email ORDER BY total_sales DESC
The expected step-by-step breakdown for the above query:
Define the main table in the FROM clause: FROM Customers AS T1.
Define the first JOIN operation: INNER JOIN.
Define the table to join: Orders AS T2.
Define the join condition: ON T1.customer_id = T2.customer_id.
Define the second JOIN operation: LEFT JOIN.
Define the table to join: OrderDetails AS T3.
Define the join condition: ON T2.order_id = T3.order_id.
Define the main filtering condition in the WHERE clause: WHERE T2.order_date >= ’2023-01-01’.
Add the additional filtering condition in the WHERE clause : AND T2.order_date <= (subquery).
Define the main table in the subquery’s FROM clause: FROM Orders.
Define the main filtering condition in the subquery’s WHERE clause: WHERE order_id = ’1’.
Select the column to be included in the subquery result: SELECT order_date.
Order the subquery results by the specified column: ORDER BY order_date DESC.
Limit the subquery results: LIMIT 1.
Complete the filtering condition in the WHERE clause: AND T2.order_date <= (SELECT order_date FROM Orders WHERE order_id = ’1’ ORDER BY order_date DESC LIMIT 1).
Group the results by the specified columns: GROUP BY T1.name, T1.email.
Select the columns to be included in the final result: SELECT T1.name, T1.email, SUM(T3.amount) AS total_sales.
Order the results by the specified column: ORDER BY total_sales DESC.
Now, I will provide you with a query, and I expect you to respond in this format:
${\{sql\_query\}}$
Figure S3. Prompt used for generating reasoning steps.
@@ A.6. Prompts
Figure S4. Example of a prompt used for the schema linking step in our proposed text-to-SQL full pipeline. The prompt is marked with yellow background, while the response is denoted with blue background. Database schemas are represented in form of DDL statements as described in Section 3.1. We used the LangChain (LangChain, [n. d.]) library to query LLMs, leveraging its Python dataclass serialization support. Note: the injected instructions for the output JSON format are not included in this example.
A.7. RetrySQL generation examples
@@
Figure S5. RetrySQL generation example with prompt and completion. The prompt is marked with yellow background, while the response is denoted with blue background. The model corrects itself after erroneously generating a reasoning step in its CoT process. Specifically, the model generates an additional filtering condition when the main one has not been declared yet. After generating the [BACK] token, it retries and correctly generates the main filtering condition, and then proceeds until the end of the reasoning chain, finally producing a correct SQL query.
@@
Figure S6. RetrySQL generation example with prompt and completion. The prompt is marked with yellow background, while the response is denoted with blue background. The model corrects itself after erroneously trying to halt the reasoning chain too early. This happens two times along the process. The retry mechanism allows the model to understand the mistakes and correct itself. Specifically, the model corrects itself by adding additional WHERE and ORDER operations. After each correction, the model proceeds to generate the rest of the reasoning chain, finally producing a correct SQL query.
@@
Figure S7. RetrySQL generation example with prompt and completion. The prompt is marked with yellow background, while the response is denoted with blue background. The model corrects itself after erroneously trying to halt the reasoning chain too early. This happens three times along the process. The retry mechanism allows the model to understand the mistakes and correct itself. In particular, the model corrects itself by adding additional JOIN and WHERE operations. After each correction, the model proceeds to generate the rest of the reasoning chain, finally producing a correct SQL query.

### v2

@@ A.7 Training data example
Figure S2: Example of a training data sample used in our experiments. It consists of the following elements: DDL statements for schema representation, external knowledge and question extracted from the BIRD metadata, reasoning steps generated as described in Section 3.2 and the ground truth SQL query. These components are separated by special tokens to guide the model in the learning process.
A.8 Prompts
You are an SQL expert. When I provide you with a SQL query, your task is to describe step-by-step how a person would create such a query. Follow the standard SQL execution order. Write the steps from the perspective of someone constructing the query. If the query includes subqueries, describe them step-by-step in the same detailed manner as the main query before referencing them. Each step should represent a distinct operation. Make the operations as granular as possible. Here is the standard SQL execution order you should follow for your explanation. If a given clause is not present in the query, skip it without mentioning its absence: 1. FROM clause (including JOINs). 2. WHERE clause. 3. GROUP BY clause. 4. HAVING clause. 5. SELECT clause. 6. ORDER BY clause. 7. LIMIT clause. For each query I provide: 1. Explain the query step by step in plain language. 2. Ensure that each step corresponds to one small, logical operation. 3. Use clear and concise language for each operation. 4. Each step should be provided in a single line (use single newline character between steps). Here is an example query for your reference: SELECT T1.name, T1.email, SUM(T3.amount) AS total_sales FROM Customers AS T1 INNER JOIN Orders AS T2 ON T1.customer_id = T2.customer_id LEFT JOIN OrderDetails AS T3 ON T2.order_id = T3.order_id WHERE T2.order_date >= ’2023-01-01’ AND T2.order_date <= (SELECT order_date FROM Orders WHERE order_id = ’1’ ORDER BY order_date DESC LIMIT 1) GROUP BY T1.name, T1.email ORDER BY total_sales DESC The expected step-by-step breakdown for the above query: Define the main table in the FROM clause: FROM Customers AS T1. Define the first JOIN operation: INNER JOIN. Define the table to join: Orders AS T2. Define the join condition: ON T1.customer_id = T2.customer_id. Define the second JOIN operation: LEFT JOIN. Define the table to join: OrderDetails AS T3. Define the join condition: ON T2.order_id = T3.order_id. Define the main filtering condition in the WHERE clause: WHERE T2.order_date >= ’2023-01-01’. Add the additional filtering condition in the WHERE clause : AND T2.order_date <= (subquery). Define the main table in the subquery’s FROM clause: FROM Orders. Define the main filtering condition in the subquery’s WHERE clause: WHERE order_id = ’1’. Select the column to be included in the subquery result: SELECT order_date. Order the subquery results by the specified column: ORDER BY order_date DESC. Limit the subquery results: LIMIT 1. Complete the filtering condition in the WHERE clause: AND T2.order_date <= (SELECT order_date FROM Orders WHERE order_id = ’1’ ORDER BY order_date DESC LIMIT 1). Group the results by the specified columns: GROUP BY T1.name, T1.email. Select the columns to be included in the final result: SELECT T1.name, T1.email, SUM(T3.amount) AS total_sales. Order the results by the specified column: ORDER BY total_sales DESC. Now, I will provide you with a query, and I expect you to respond in this format: $\{sql\_query\}$}
Figure S3: Prompt used for generating reasoning steps.
@@ A.8 Prompts
Figure S4: Example of a prompt used for the schema linking step in our proposed text-to-SQL full pipeline. The prompt is marked with yellow background, while the response is denoted with blue background. Database schemas are represented in form of DDL statements as described in Section 3.1. We used the LangChain library to query LLMs, leveraging its Python dataclass serialization support. Note: the injected instructions for the output JSON format are not included in this example.
A.9 RetrySQL generation examples
@@
Figure S5: RetrySQL generation example with prompt and completion. The prompt is marked with yellow background, while the response is denoted with blue background. The model corrects itself after erroneously generating a reasoning step in its CoT process. Specifically, the model generates an additional filtering condition when the main one has not been declared yet. After generating the [BACK] token, it retries and correctly generates the main filtering condition, and then proceeds until the end of the reasoning chain, finally producing a correct SQL query.
@@
Figure S6: RetrySQL generation example with prompt and completion. The prompt is marked with yellow background, while the response is denoted with blue background. The model corrects itself after erroneously trying to halt the reasoning chain too early. This happens two times along the process. The retry mechanism allows the model to understand the mistakes and correct itself. Specifically, the model corrects itself by adding additional WHERE and ORDER operations. After each correction, the model proceeds to generate the rest of the reasoning chain, finally producing a correct SQL query.
@@
Figure S7: RetrySQL generation example with prompt and completion. The prompt is marked with yellow background, while the response is denoted with blue background. The model corrects itself after erroneously trying to halt the reasoning chain too early. This happens three times along the process. The retry mechanism allows the model to understand the mistakes and correct itself. In particular, the model corrects itself by adding additional JOIN and WHERE operations. After each correction, the model proceeds to generate the rest of the reasoning chain, finally producing a correct SQL query.
– External Knowledge: unknown power refers to power = ’*’ or POWER IS NULL; commander play format refers to format = ’commander’; legal for commander play format refers to format = ’commander’ where status = ’Legal’ Define the main table in the FROM clause: ‘FROM cards AS T1‘. Define the JOIN operation: ‘INNER JOIN‘. Define the table to join: ‘legalities AS T2‘. Define the join condition: ‘ON T1.uuid = T2.uuid‘. Define the main filtering condition in the WHERE clause: ‘WHERE T1.power = ’*’‘. Add the additional filtering condition in the WHERE clause: ‘AND T2.format = ’commander’‘. Add the additional filtering condition in the WHERE clause: ‘AND T2.status = ’Legal’‘. Select the columns to be included in the final result: ‘SELECT T1.id, T1.artist‘. [SQL]
SELECT T1.id, T1.artist FROM cards AS T1 INNER JOIN legalities AS T2 ON T1.uuid = T2.uuid WHERE T1.power = ’*’ AND T2.format = ’commander’ AND T2.status = ’Legal’
Define the main table in the FROM clause: ‘FROM cards AS T1‘. Define the JOIN operation: ‘INNER JOIN‘. Define the table to join: ‘legalities AS T2‘. Define the join condition: ‘ON T1.uuid = T2.uuid‘. Define the main filtering condition in the WHERE clause: ‘WHERE T1.power = ’*’‘. Select the columns to be included in the final result: ‘SELECT T1.id, T1.artist‘. [BACK]
Add the additional filtering condition in the WHERE clause: ‘OR T1.power IS NULL‘.
Select the columns to be included in the final result: ‘SELECT T1.id, T1.artist‘.
[BACK]
Add the additional filtering condition in the WHERE clause: ‘AND T2.status = ’Legal’‘.
Select the columns to be included in the final result: ‘SELECT T1.id, T1.artist‘. [SQL]
SELECT T1.id, T1.artist FROM cards AS T1 INNER JOIN legalities AS T2 ON T1.uuid = T2.uuid WHERE T1.power = ’*’ OR T1.power IS NULL AND T2.format = ’commander’ AND T2.status = ’Legal’
Figure S8: Comparison of error-free and RetrySQL model generation. The external evidence is marked with yellow background, error-free response has a pink background, while the RetrySQL response is denoted with blue background. The error-free model omits one condition in the external knowledge (POWER IS NULL). The RetrySQL-trained model utilizes the retry mechanism to go back from an incorrect step and generate the full condition.
– External Knowledge: triple bond refers to bond_type = ’#’; phosphorus refers to element = ’p’; bromine refers to element = ’br’ Define the main table in the FROM clause: ‘FROM atom AS T1‘. Define the first JOIN operation: ‘INNER JOIN‘. Define the table to join: ‘bond AS T2‘. Define the join condition: ‘ON T1.molecule_id = T2.molecule_id‘. Define the second JOIN operation: ‘INNER JOIN‘. Define the table to join: ‘molecule AS T3‘.
Define the join condition: ‘ON T2.molecule_id = T3.molecule_id‘.
Define the main filtering condition in the WHERE clause: ‘WHERE T2.bond_type = ’triple’‘.
Add the additional filtering condition in the WHERE clause: ‘AND (T1.element = ’p’ OR T1.element = ’br’)‘.
Select the column to be included in the final result: ‘SELECT COUNT(T1.atom_id)‘. [SQL]
SELECT COUNT(T1.atom_id) FROM atom AS T1 INNER JOIN bond AS T2 ON T1.molecule_id = T2.molecule_id INNER JOIN molecule AS T3 ON T2.molecule_id = T3.molecule_id WHERE T2.bond_type = ’triple’ AND (T1.element = ’p’ OR T1.element = ’br’)
Define the main table in the FROM clause: ‘FROM Molecule AS T1‘. Define the first JOIN operation: ‘INNER JOIN‘. Define the table to join: ‘Atom AS T2‘. Define the join condition: ‘ON T1.molecule_id = T2.molecule_id‘. Define the second JOIN operation: ‘INNER JOIN‘. Define the table to join: ‘Bond AS T3‘. Define the join condition: ‘ON T1.molecule_id = T3.molecule_id‘. Define the main filtering condition in the WHERE clause: ‘WHERE T2.element IN (’p’, ’br’)‘. Select the column to be included in the final result: ‘SELECT COUNT(T2.atom_id)‘. [BACK]
Add the additional filtering condition in the WHERE clause: ‘AND T3.bond_type = ’#’‘.
Select the column to be included in the final result: ‘SELECT COUNT(T2.atom_id)‘.
[SQL]
SELECT COUNT(T2.atom_id) FROM Molecule AS T1 INNER JOIN Atom AS T2 ON T1.molecule_id = T2.molecule_id INNER JOIN Bond AS T3 ON T1.molecule_id = T3.molecule_id WHERE T2.element IN (’p’, ’br’) AND T3.bond_type = ’#’
Figure S9: Comparison of error-free and RetrySQL model generation. The external evidence is marked with yellow background, error-free response has a pink background, while the RetrySQL response is denoted with blue background. The error-free model hallucinates bond_type = ‘triple‘. The RetrySQL-trained model utilizes the retry mechanism to go back from an incorrect step and generate the correct condition: bond_type = ‘#‘.
– External Knowledge: known as Archenemy refers to mcmName = ’Archenemy’; having the code ARC refers to setCode = ’ARC’ Define the main table in the FROM clause: ‘FROM sets AS T1‘. Define the JOIN operation: ‘INNER JOIN‘. Define the table to join: ‘set_translations AS T2‘. Define the join condition: ‘ON T1.id = T2.id‘. Define the main filtering condition in the WHERE clause: ‘WHERE T1.mcmName = ’Archenemy’‘. Add the additional filtering condition in the WHERE clause: ‘AND T1.code = ’ARC’‘. Select the column to be included in the final result: ‘SELECT T2.language‘. [SQL]
SELECT T2.language FROM sets AS T1 INNER JOIN set_translations AS T2 ON T1.id = T2.id WHERE T1.mcmName = ’Archenemy’ AND T1.code = ’ARC’
Define the main table in the FROM clause: ‘FROM sets AS T1‘. Define the JOIN operation: ‘INNER JOIN‘. Define the table to join: ‘set_translations AS T2‘. Define the join condition: ‘ON T1.code = T2.setcode‘.
Define the main filtering condition in the WHERE clause: ‘WHERE T1.mcmName = ’Archenemy’‘.
Select the column to be included in the final result: ‘SELECT T2.language‘. [BACK]
Add the additional filtering condition in the WHERE clause: ‘AND T2.setcode = ’ARC’‘.
Select the column to be included in the final result: ‘SELECT T2.language‘. [SQL]
SELECT T2.language FROM sets AS T1 INNER JOIN set_translations AS T2 ON T1.code = T2.setcode WHERE T1.mcmName = ’Archenemy’ AND T2.setcode = ’ARC’
Figure S10: Comparison of error-free and RetrySQL model generation. The external evidence is marked with yellow background, error-free response has a pink background, while the RetrySQL response is denoted with blue background. The error-free model hallucinates the join condition ON T1.id = T2.id. The RetrySQL-trained model generates the correct condition: ON T1.code = T2.setcode.
Define the main table in the FROM clause: ‘FROM superhero AS T1‘. Define the first JOIN operation: ‘INNER JOIN‘. Define the table to join: ‘colour AS T2‘. Define the join condition: ‘ON T1.eye_colour_id = T2.id‘. Define the second JOIN operation: ‘INNER JOIN‘. Define the table to join: ‘colour AS T3‘. Define the join condition: ‘ON T1.hair_colour_id = T3.id‘. Define the third JOIN operation: ‘INNER JOIN‘. Define the table to join: ‘colour AS T4‘. Define the join condition: ‘ON T1.skin_colour_id = T4.id‘. Define the fourth JOIN operation: ‘INNER JOIN‘. Define the table to join: ‘gender AS T5‘. Define the join condition: ‘ON T1.gender_id = T5.id‘. Define the fifth JOIN operation: ‘INNER JOIN‘. Define the table to join: ‘race AS T6‘. Define the join condition: ‘ON T1.race_id = T6.id‘. Define the sixth JOIN operation: ‘INNER JOIN‘. Define the table to join: ‘publisher AS T7‘. Define the join condition: ‘ON T1.publisher_id = T7.id‘. Define the seventh JOIN operation: ‘INNER JOIN‘. Define the table to join: ‘alignment AS T8‘. Define the join condition: ‘ON T1.alignment_id = T8.id‘. Define the main filtering condition in the WHERE clause: ‘WHERE T2.colour = ’Blue’‘. Add the additional filtering condition in the WHERE clause: ‘AND T1.eye_colour_id = T2.id‘. Add the additional filtering condition in the WHERE clause: ‘AND T1.hair_colour_id = T3.id‘. Add the additional filtering condition in the WHERE clause: ‘AND T1.skin_colour_id = T4.id‘. Add the additional filtering condition in the WHERE clause: ‘AND T1.gender_id = T5.id‘. Add the additional filtering condition in the WHERE clause: ‘AND T1.race_id = T6.id‘. Add the additional filtering condition in the WHERE clause: ‘AND T1.publisher_id = T7.id‘. Add the additional filtering condition in the WHERE clause: ‘AND T1.alignment_id = T8.id‘. Select the column to be included in the final result: ‘SELECT COUNT(T1.id)‘. [SQL]
SELECT COUNT(T1.id) FROM superhero AS T1 INNER JOIN colour AS T2 ON T1.eye_colour_id = T2.id INNER JOIN colour AS T3 ON T1.hair_colour_id = T3.id INNER JOIN colour AS T4 ON T1.skin_colour_id = T4.id INNER JOIN gender AS T5 ON T1.gender_id = T5.id INNER JOIN race AS T6 ON T1.race_id = T6.id INNER JOIN publisher AS T7 ON T1.publisher_id = T7.id INNER JOIN alignment AS T8 ON T1.alignment_id = T8.id WHERE T2.colour = ’Blue’ AND T1.eye_colour_id = T2.id AND T1.hair_colour_id = T3.id AND T1.skin_colour_id = T4.id AND T1.gender_id = T5.id AND T1.race_id = T6.id AND T1.publisher_id = T7.id AND T1.alignment_id = T8.id
Define the main table in the FROM clause: ‘FROM superhero AS T1‘. Define the JOIN operation: ‘INNER JOIN‘. Define the table to join: ‘colour AS T2‘. Define the join condition: ‘ON T1.eye_colour_id = T2.id‘. Define the main filtering condition in the WHERE clause: ‘WHERE T2.colour = ’Blue’‘. Select the column to be included in the final result: ‘SELECT COUNT(T1.id)‘. [BACK]
Add the additional filtering condition in the WHERE clause: ‘AND T1.eye_colour_id = T2.id‘.
Select the column to be included in the final result: ‘SELECT COUNT(T1.id)‘. [SQL]
SELECT COUNT(T1.id) FROM superhero AS T1 INNER JOIN colour AS T2 ON T1.eye_colour_id = T2.id WHERE T2.colour = ’Blue’ AND T1.eye_colour_id = T2.id
Figure S11: Comparison of error-free and RetrySQL model generation. The error-free response has a pink background, while the RetrySQL response is denoted with blue background. The error-free model generates many more reasoning steps than the RetrySQL-trained model - it does not have the ability to backtrack, so it can only generate more steps to solve a given task.